Assume that when adults with smartphones are randomly​ selected, 45​% use them in meetings or classes. If 5 adult smartphone users are randomly​ selected, find the probability that at least 3 of them use their smartphones in meetings or classes.
Answer by Boreal(15235)

This is 3,4,5
for 5 it is 0.45^5=0.0185
for 4 it is 5(0.45)^4*0.55=0.1128; the 5 is the number of ways it can occur.
for 3 it is 5C3*0.43^3*0.55^2=0.2757
that sum is 0.4070
Can do 1-binomcdf(5,.45,2 )on the calculator and that removes the 0,1,2 probability from the distribution.